# sEMG-Driven Physics-Informed Gated Recurrent Networks for Modeling Upper Limb Multi-Joint Movement Dynamics

### v1

@@ Abstract
Exoskeletons and rehabilitation systems offer great potential for enhancing human strength and recovery through advanced human-machine interfaces (HMIs) that adapt to movement dynamics. However, the real-time application of physics-informed neural networks (PINNs) is limited by their reliance on fixed input lengths and surrogate models. This study introduces a novel physics-informed Gated Recurrent Network (PiGRN) designed to predict multi-joint torques using surface electromyography (sEMG) data. The PiGRN model employs a Gated Recurrent Unit (GRU) to convert time-series sEMG inputs into multi-joint kinematics and external loads, which are then integrated into an equation of motion to ensure consistency with physical laws. Experimental validation with sEMG data from five participants performing elbow flexion-extension tasks showed that the PiGRN model accurately predicted joint torques for 10 unfamiliar movements, with RMSE values between 4.02% and 11.40% and correlation coefficients ranging from 0.87 to 0.98. These findings highlight the PiGRN’s potential for real-time exoskeleton and rehabilitation applications. Future research will explore more diverse datasets, improve musculoskeletal models, and investigate unsupervised learning methods.
@@ 1 Introduction
In recent years, the Gated Recurrent Unit (GRU), introduced by Kyunghyun Cho et al. in 2014 [36], offers a simplified alternative to traditional RNNs, resembling LSTM networks [37]. GRUs streamline structure reduces parameter count while maintaining performance. They are proficient in analyzing time series data, GRUs excel in capturing long dependencies without vanishing gradient issues [38, 39]. Furthermore, their recursive design enables them to sustain and update a hidden state across time, efficiently handling sequential data and capturing temporal patterns [40]. Hence GRUs are ideal for processing sEMG signal time series, representing muscle activity patterns effectively.
This work aims to achieve three objectives:
@@
    Analysis of sEMG for Predicting Joint Torque: We propose a novel physics-informed Gated Recurrent Network (PiGRN) to analyze sEMG signals to predict joint torques. Also, the GRU within the PiGRN extracts temporal features from sEMG time series, accommodating variable input sequences.
@@
    Correlating sEMG with Joint Kinematics and External Load: In the proposed framework the GRU component first maps sEMG signals to joint kinematic parameters (angle, velocity, acceleration) and external load, while adhering with dynamic motion equations. This integration serves as a soft constraint on predicted kinematics and load.
@@ 3. III
    Validation Across Multiple Degrees of Freedom and Load: We validated the robustness of PiGRN model in predicting joint torques across multiple degrees of freedom, considering varied muscle loading scenarios in an upper limb model. This includes accurately predicting shoulder and elbow joint torques during elbow FE tasks using sEMG recordings from specified muscles under different loads.
@@ 6 CONCLUSION
This paper presents a novel physics-informed gated recurrent neural network (PiGRN) developed to predict joint torque using sEMG data under various load conditions. The proposed method is capable of predicting multi-joint kinematic parameters and external loads from sEMG sequences of different lengths. Unlike conventional deep learning approaches, our framework uniquely integrates predicted joint kinematic parameters and external loads into physics-based domain knowledge as soft constraints to regularize the GRU’s loss function, a technique that has not been implemented before. This strategy reduces the dependence on automatic differentiation and decreases training time. Experiments with sEMG data across three different load groups demonstrate the method’s effectiveness in predicting joint torque and kinematics. We believe this framework could serve as a general method for predicting joint torque and kinematics, making it well-suited for real-time applications in the development of robust HMIs. Although this study focused on data involving two degrees of freedom, two joints, and in-plane movements, future work will aim to address more generalized datasets and musculoskeletal models, as well as explore unsupervised learning techniques.

### v2

@@ Abstract
Exoskeletons and rehabilitation systems have the potential to improve human strength and recovery by using adaptive human-machine interfaces. Achieving precise and responsive control in these systems depends on accurately estimating joint movement dynamics, such as joint angle, velocity, acceleration, external mass, and torque. While machine learning (ML) approaches have been employed to predict joint kinematics from surface electromyography (sEMG) data, traditional ML models often struggle to generalize across dynamic movements. In contrast, physics-informed neural networks integrate biomechanical principles, but their effectiveness in predicting full movement dynamics has not been thoroughly explored. To address this, we introduce the Physics-informed Gated Recurrent Network (PiGRN), a novel model designed to predict multi-joint movement dynamics from sEMG data. PiGRN uses a Gated Recurrent Unit (GRU) to process time-series sEMG inputs, estimate multi-joint kinematics and external loads, and predict joint torque while incorporating physics-based constraints during training. Experimental validation, using sEMG data from five participants performing elbow flexion-extension tasks with 0 kg, 2 kg, and 4 kg loads, showed that PiGRN accurately predicted joint torques for 10 novel movements. RMSE values ranged from 4.02% to 11.40%, with correlation coefficients between 0.87 and 0.98. These results underscore PiGRN’s potential for real-time applications in exoskeletons and rehabilitation. Future work will focus on expanding datasets, improving musculoskeletal models, and investigating unsupervised learning approaches.
@@ 1 Introduction
In recent years, the Gated Recurrent Unit (GRU), introduced by Kyunghyun Cho et al. in 2014 [36], offers a simplified alternative to traditional RNNs, resembling LSTM networks [37]. GRUs streamline structure reduces parameter count while maintaining performance. They are proficient in analyzing time series data, GRUs excel in capturing long dependencies without vanishing gradient issues [38, 39]. Furthermore, their recursive design enables them to sustain and update a hidden state across time, efficiently handling sequential data and capturing temporal patterns [40]. Hence GRUs are ideal for processing sEMG signal time series, representing muscle activity patterns effectively. This work aims to achieve the following objectives:
@@
    Predicting joint dynamics using a Physics-Informed Gated Recurrent Network (PiGRN) model: We developed a PiGRN model by integrating a Gated Recurrent Unit (GRU) with physics-based constraints to process time-series sEMG data and predict multi-joint movement dynamics parameters, such as joint angle, velocity, acceleration, and external mass. Which is further used to estimate joint torque.
@@
    Experimental validation of PiGRN for joint dynamics prediction: To assess the performance of PiGRN, we have collected sEMG data from five participants performing elbow flexion-extension tasks under varying external loads (0 kg, 2 kg, and 4 kg).
@@ 3. III
    Generalization across multiple joints and load: We validated the robustness of the PiGRN model in predicting multi-joint dynamics across varied muscle loading scenarios in an upper limb model.
@@ 6 CONCLUSION
This paper presents a novel physics-informed gated recurrent neural network (PiGRN) developed to predict multi-joint dynamics using sEMG data under various load conditions. The proposed method is capable of predicting multi-joint kinematic parameters and external loads from sEMG sequences of different lengths. Unlike conventional deep learning approaches, our framework uniquely integrates predicted joint kinematic parameters and external loads into physics-based domain knowledge as soft constraints to regularize the GRU’s loss function, a technique that has not been implemented before. This strategy reduces the dependence on automatic differentiation and decreases training time. Experiments with sEMG data across three different load groups demonstrate the method’s effectiveness in predicting joint torque and kinematics. We believe this framework could serve as a general method for predicting joint torque and kinematics, making it well-suited for real-time applications in the development of robust HMIs. Although this study focused on data involving two degrees of freedom, two joints, and in-plane movements, future work will aim to address more generalized datasets and musculoskeletal models, as well as explore unsupervised learning techniques.